# Prot2Chat: Protein LLM with Early Fusion of Sequence and Structure

### v1

@@ Abstract
Proteins play a pivotal role in living organisms, yet understanding their functions presents significant challenges, including the limited flexibility of classification-based methods, the inability to effectively leverage spatial structural information, and the lack of systematic evaluation metrics for protein Q&A systems. To address these limitations, we propose Prot2Chat, a novel framework that integrates multimodal protein representations with natural language through a unified module, enabling large language model (LLM)-driven answer generation. Our model incorporates a modified ProteinMPNN encoder, which encodes protein sequence and structural information in a unified manner, a protein-text adapter with cross-attention mechanisms, and a LLaMA3 decoder. To optimize training efficiency, we freeze the encoder and employ LoRA techniques for the decoder. We conducted experiments on two datasets, both automated metrics and expert evaluations demonstrate the superior performance of our model. Furthermore, zero-shot prediction results highlight its strong generalization capabilities. This framework offers a promising solution for bridging protein domain knowledge with natural language understanding, paving the way for transformative advancements in protein-related research.
@@ 1 Introduction
With the continuous advancement of biotechnology, gaining a deeper understanding of protein sequences, structures, and their functions has become increasingly important. To elucidate the complex mechanisms underlying protein-mediated physiological processes, researchers have focused on studying and predicting the physicochemical properties of proteins. The widespread availability of protein sequences and structures, enabled by novel high-throughput sequencing technologies and advanced structure prediction tools Jumper et al. (2021); Abramson et al. (2024), has greatly facilitated the exploration of intricate relationships between protein sequences, structures, and functions. In this context, protein Q&A systems have emerged as a research hotspot, aiming to help scientists and researchers efficiently access information about protein properties and functions Zhou et al. (2023). These systems are of considerable significance in fields such as drug discovery and disease research. Besides, just as some proteins mentioned in AF-Cluster Wayment-Steele et al. (2024), it indicates that even if the protein sequences are the same, they may present different 3D structures due to different environmental conditions, binding partners, or point mutations, thereby performing different biological functions.
Traditional approaches, such as those based on Gene Ontology (GO) classifications, predict protein functions by leveraging sequence information through convolutional models or sequence similarity-based computations, as exemplified by methods like DeepGOPlus Kulmanov and Hoehndorf (2020). Such models primarily utilize protein sequence information and are limited to classification tasks, which oversimplifies the complexity of protein functions. The recent rise of large language models (LLMs) has propelled protein Q&A systems into a cutting-edge research domain, allowing scientists to efficiently access and interpret critical information about protein properties and functions Zhou et al. (2023). These systems hold immense potential for transformative applications in areas such as drug discovery and disease research. To support this growing field, several datasets, including Mol-Instructions Fang et al. (2023) and UniProtQA Luo et al. (2023), have been developed, providing essential resources for advancing protein Q&A research.
An increasing number of models have fine-tuned LLMs using protein-specific Q&A datasets to generate contextually relevant text or answer questions related to protein structure and function. However, as highlighted in Table 1, existing approaches often face significant limitations, including a lack of multimodal information integration, limited flexibility, and high computational costs. These shortcomings restrict their ability to provide precise and contextually appropriate responses. For instance, FAPM Xiang et al. (2024) and InstructProtein Wang et al. (2023), despite incorporating generative Q&A functionalities, remain heavily influenced by classification-based frameworks. This often results in the generation of extraneous or irrelevant information, deviating from the intended scope of the answers and limiting their utility in addressing complex protein-related queries. BiomedGPT Luo et al. (2023) introduces a generative framework tailored for protein Q&A tasks but relies solely on sequence information, which significantly limits its predictive capabilities.
We compared the generation effects of large language models trained with protein structures and sequences, respectively, on different types of questions, as shown in Figure 1. The results show that when answering questions related to protein subcellular localization and domains or motifs, the generation effect of the model jointly trained with protein structures and sequences is significantly better than that trained only with protein sequences. Some following works such as ProtChatGPT Wang et al. (2024) and Evola Zhou et al. (2025) have tried to employ two separate modules for encoding sequence and structural information. However, such practice makes it difficult for protein sequence and structural features to interact, and significantly increases computational complexity. Moreover, no existing model has systematically examined the relative contributions of sequence and structural information to the performance of protein Q&A systems. This gap underscores the critical need for future research to evaluate how integrating both sequence and structural data can enhance the accuracy, robustness, and interpretability of protein-related Q&A.
@@
As the sequence and structural features of proteins are one-to-one correspondence, we propose a novel framework called Prot2Chat that seamlessly incorporates the spatial structure and sequence information into LLMs. This approach seeks to enhance both the performance and applicability of protein Q&A systems. As illustrated in Figure 2, the overall architecture of our model comprises three primary components: the protein encoder, the protein-to-text adapter, and the large language model. Specifically, we extended the structure encoder ProteinMPNN Dauparas et al. (2022) to fuse sequence information during node initialization. The sequence features we introduced are inherent to the model, thus eliminating the need for additional training parameters. Then following the methodology of InstructPLM Qiu et al. (2024), a lightweight cross-attention layer was incorporated to align the protein encoder with a LoRA Hu et al. (2022) fine-tuned LLM. This approach preserves the primary sequence and structural information of proteins during the compression process. The processed protein structure and sequence data are provided as a soft prompt to the LLM. Leveraging the flexible text generation capabilities of the LLM, our model effectively follows residue-level protein instructions, significantly enhancing its utility in protein-related tasks.
@@
To evaluate the model’s performance, we conducted comparative experiments on the Mol-Instructions and UniprotQA datasets. And we conducted zero-shot experiments using a part of the UniprotQA dataset to verify the generalization ability of the model. The generated answers were assessed using common metrics such as BLEU and ROUGE, as well as online KIMI Qin et al. (2024) scoring metrics and expert evaluations, ensuring the reliability and relevance of the outputs. Experimental results demonstrate that the model employing a unified encoder to integrate protein sequence and structural information outperforms those using sequence-only pretraining or ESM-based encoders Lin et al. (2022). Furthermore, the high consistency between expert evaluations and online KIMI assessments confirms the robustness of our model.
@@
    We extended the existing structure encoder ProteinMPNN to realize structure and sequence early fusion without the need for training.
@@
    Based on the protein encoder, we achieved a lightweight and effective protein large language model with only 93M training parameters.
@@
Figure 1: Comparison of Q&A results between Prot2Chat and LLaMA3-FT with sequence.
Figure 2: Model Structure of Prot2Chat. The red font represents the input, the snowflake represents freezing, and the fire represents the parameters to be trained.
@@ 2 Related work
Protein Representation Learning
@@
Proteins are fundamental components of cells, essential for their biological activities and diverse functions. Previous studies on protein characterization have explored various methods to learn protein representations based on different forms of protein information. Protein sequences, often referred to as the ”language of life,” have been extensively studied using advanced natural language processing techniques. For instance, Tranception Notin et al. (2022) employs the Transformer Vaswani (2017) model to encode amino acid sequences, capturing relationships between residues and autoregressively reconstructing protein sequences from large-scale databases. Similarly, sequential modeling approaches such as ESM Rives et al. (2021); Lin et al. (2023); Hayes et al. (2024) leverage masked language modeling (MLM) to develop attention patterns that correspond to residue-residue contact maps, enhancing sequence-based protein representations. On the other hand, structure-based approaches Gligorijević et al. (2021); Achiam et al. (2023) directly indicates protein function and encodes geometric information of the protein for topology-sensitive tasks such as protein property prediction. Foldseek van Kempen et al. (2022) introduces the concept of a structural alphabet, encoding protein structures into a discrete representation space. Similarly, Saprot Su et al. (2023) introduces a structure-aware vocabulary, embedding structural information into model inputs to enhance representational capacity, achieving significant success in protein function prediction tasks.
Multi-Modal Alignment
@@
Enabling large language models (LLMs) to understand additional modalities has been a rapidly evolving research direction, with notable examples including image-text models Tsimpoukelli et al. (2021); Li et al. (2023), video-text models like VideoLlama Zhang et al. (2023), audio-text models such as Macaw-LLM Lyu et al. (2023), and molecular-text models, including ReLM Shi et al. (2023), and MolTC Fang et al. (2024). This line of research was pioneered by advancements in visual language modeling (VLM), which has been successfully applied to tasks such as few-shot image classification, image captioning, and image Q&A Li et al. (2023); Alayrac et al. (2022). To enable LLMs to understand images, leading VLM methods adopt different strategies. Some, like BLIP-2 Li et al. (2023), employ nonlinear and expressive cross-modal projectors, while others, such as PaLM-E Driess et al. (2023), utilize visual encoders and fine-tune LLMs on multimodal datasets. These approaches have also been increasingly applied to protein analysis and understanding. For instance, Galactica Taylor et al. (2022) leverages scientific literature to model protein sequences and SMILES representations, enabling the model to interpret sequence properties. ProtNote Char et al. (2024), utilize two encoders to encode protein sequence information and text information, respectively, and Multilayer Perceptron (MLP) to fuse these inputs, achieving the goal of using text to achieve supervised and zero-shot protein functional prediction. ProteinChat Guo et al. (2023) further uses the protein structure and the corresponding description to model for the protein Q&A tasks. ProtChatGPT Wang et al. (2024) uses two independent modules to encode sequence and structural information respectively, and aligns them with natural language to explore protein functions. InstructBioMol Zhuang et al. (2024) integrates multimodal biomolecular inputs through a feature extraction module, allowing researchers to articulate design goals in natural language. InstructProtein Wang et al. (2023) employs a supervised training framework based on instruction generation using knowledge graphs, enabling bidirectional generation between natural language and protein language. This model can predict protein functional descriptions and generate protein sequences based on natural language prompts. Prot2Text Abdine et al. (2024) combines graph neural networks (GNNs) with LLMs to predict protein functions in free-text form. Similarly, ProtT3 Liu et al. (2024) provides an efficient protein-to-text generation framework by integrating protein language models (PLMs) with LLMs. FAPM Xiang et al. (2024) utilizes a contrastive learning framework to implement a generative-like Q&A model, while BiomedGPT Luo et al. (2023) bridges the gap between biological and human natural language through a multimodal generative pre-trained transformer (GPT). This allows users to ”communicate” with biological modalities using free-text queries, facilitating interactions with biological data in natural language.
@@
Prot2Chat [Structure+Sequence]&Question Free Text
Table 1: Comparison of input-output of protein Q&A models, ”Description” is a text related to protein functions, such as the description of Gene Ontology (GO) terms, ”[Structure + Sequence]” represents the early fusion of structure and sequence.
@@
Table 2: The count and division of the datasets we used.
Prompt
Please select the sentence that is closest and most accurate in meaning to the Target sentence and rank sentences A, B, and C accordingly, with the first place meaning the most identical and accurate in meaning to the Target. Only provide the answer, for example, ‘B A C’ or ‘A C B’, etc.
Target: {Ground Truth} A: {Model1 Generated Text} B: {Model2 Generated Text} C: {Model3 Generated Text}
Table 3: Prompt to KIMI for evaluation
@@
Prot2Chat 33.25 54.88 35.26 47.90
@@
Prot2Chat (zero-shot) 2.11 13.78 1.51 9.64
Prot2Chat (fine-tuned) 5.87 15.65 9.00 14.99
Model 1st 2nd 3rd Avg.
Prot2Chat 591 386 22 1.43
BioMedGPT-LM-10B 218 214 567 2.35
LLaMA3-FT with sequence 190 399 410 2.22
@@
Model 1st 2nd 3rd Avg.
Prot2Chat 454 115 78 1.42
BioMedGPT-LM-10B 157 203 287 2.20
LLaMA3-FT with sequence 36 329 282 2.38
@@
[Uncaptioned image] [O58787]Using the provided structure and amino acid sequence of a protein, estimate its primary location within the cell: Based on the provided amino acid sequence, it appears that the protein is a membrane protein, likely localized to the endoplasmic reticulum, Golgi apparatus, and/or plasma membrane. The presence of multiple transmembrane domains and the distribution of hydrophobic and hydrophilic residues suggest that this protein is an integral membrane protein. Additionally, the sequence contains several potential phosphorylation sites and potential N-glycosylation sites, which may be involved in protein-protein interactions and/or post-translational modification. The protein characterized by the structure and amino acid sequence demonstrates rRNA binding and is implicated in the translation. Its subcellular localization is primarily within the cytosol. Upon analyzing the given protein structure and sequence, it appears that the protein is likely to be localized in the cytosol compartment of the cell.
[Uncaptioned image] [O58736]Using the protein structure and the following protein sequence as input, assess its potential function by comparing it to known proteins with similar sequences or structures: The provided protein sequence is approximately 240 amino acids long and shows no significant similarity to known proteins in the database. However, it can be analyzed for potential functional regions and motifs using various bioinformatic tools. Here are some observations and predictions: 1. **Transmembrane regions**: The sequence contains several regions that are predicted to be transmembrane, including the following: * 24-45 amino acids (ILLLVGMPGSGKGEVAKAFRRRGIPVI) * 63-82 amino acids (MGGDAIREEAEKRGIPKTPEGLKYVSLKVREE) With the provided structure and amino acid sequence, the protein is expected to display adenylate kinase activity and participate in the phosphorylation, ATP biosynthetic process, phosphorylation. The protein with the structure and amino acid sequence has been analyzed, and its structure suggests a potential involvement in ATP binding.

### v2

@@ Abstract
Motivation: Proteins are of great significance in living organisms. However, understanding their functions encounters numerous challenges, such as insufficient integration of multimodal information, a large number of training parameters, limited flexibility of classification-based methods, and the lack of systematic evaluation metrics for protein Q&A systems. To tackle these issues, we propose the Prot2Chat framework.
Results: We modified ProteinMPNN to encode protein sequence and structural information in a unified way. We used a large language model (LLM) to encode questions into vectors and developed a protein-text adapter to compress protein information into virtual tokens based on these vectors, achieving the early fusion of text and protein information. Finally, the same LLM reads the virtual tokens and the questions to generate answers. To optimize training efficiency, we froze the encoder and employed Low-Rank Adaptation (LoRA) techniques for the LLM. Experiments on two datasets show that both automated metrics and expert evaluations demonstrate the superior performance of our model, and zero-shot prediction results highlight its generalization ability. The models and codes are available at https://github.com/wangzc1233/Prot2Chat.
Contact: zqcao@suda.edu.cn or wangzc025@163.com
keywords:
Protein Q&A, Early-Fusion, LLM
@@ 1 Introduction
With the continued advancement of biotechnology, gaining a deeper understanding of protein sequences, structures, and their functions has become increasingly critical. Protein function is fundamentally governed by its amino acid sequence, which encodes the structural and biochemical features necessary for activity. However, even identical protein sequences can adopt distinct three-dimensional conformations under varying environmental conditions or subtle mutations, thereby exhibiting diverse biological functions, as exemplified by cases reported in AFcluster wayment2024predicting . The widespread availability of protein sequence and structure data—enabled by high-throughput sequencing technologies and powerful structure prediction tools abramson2024accurate has significantly advanced our understanding of the physicochemical properties of proteins. In parallel, significant research efforts have focused on investigating and predicting protein functions, where deep learning models have employed sequence or structure representations as input to predict function-related properties such as Gene Ontology (GO) terms kulmanov2020deepgoplus ; su2023saprot . The recent emergence of large language models (LLMs) has propelled protein question answering (Q&A) systems at the forefront of computational biology research zhou2023survey . The related systems enable researchers to efficiently access and interpret key information regarding protein properties and functions. This holds great potential for transformative applications in fields such as drug discovery and disease research smith2025funcfetch . To support this growing field, several datasets, such as Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt , have been introduced, providing essential resources for advancing protein Q&A systems. An increasing number of models have fine-tuned LLMs on protein-specific Q&A datasets to generate contextually relevant text or to answer questions related to protein structure and function zhou2025decoding .
However, as summarized in Table 1, existing methods face several critical limitations, including inadequate integration of multimodal information, restricted flexibility, high training parameter volume, and a lack of systematic evaluation metrics. These challenges constrain the models’ ability to generate accurate and context-aware responses. For instance, although models such as FAPM xiang2024fapm and InstructProtein wang2023instructprotein incorporate generative Q&A capabilities, they remain heavily dependent on classification-based paradigms. This often leads to the production of unexpected content, deviating from the intended scope of inquiry and reducing the models’ effectiveness in handling complex protein-related questions. BiomedGPT luo2023biomedgpt enables users to interact with biological data through natural language. Nonetheless, it relies exclusively on sequence-level information, which substantially limits its predictive capacity. ProtChatGPT wang2024protchatgpt and Evola zhou2025decoding attempt to address this limitation by employing separate modules for encoding sequence and structural information. However, this modular design hinders effective interaction between sequence and structural features and significantly increases computational complexity. Moreover, no existing model has systematically examined the relative contributions of sequence and structural information to the performance of protein Q&A systems.
Figure 1: Prot2Chat can assist human in understanding protein information and achieve cross-modal information communication. Among them, ‘$<$Soft Prompt$>$‘ is the prompt obtained by our model through fusing protein structure, sequence, and text information, which helps the LLM generate more valuable answers.
@@
As the sequence and structural features of proteins are in inseparable correspondence, both jointly determine the properties of proteins. We propose a novel framework named Prot2Chat. This framework can seamlessly integrate the spatial structure and sequence information of proteins into LLMs. Meanwhile, by incorporating text information during the processing of protein information, the LLM is enabled to understand the question, so as to achieve question-based compression of protein information, and ultimately realize the early fusion of text, structure, and sequence. This approach seeks to enhance both the performance and applicability of protein Q&A systems. To be specific, we introduce protein sequence information during the initialization of nodes features in ProteinMPNN dauparas2022robust and early-fuse the protein sequence information in the protein feature representation. Then inspired by BLIP-2 li2023blip , an adapter module was incorporated to align the protein encoder with a LoRA hu2022lora fine-tuned LLM. In addition, we use the LLM to encode the question text and add it to the adapter, integrating it into the above-mentioned compression process to further achieve the early-fusion of protein and text information, aiming to improve the accuracy and robustness of protein Q&A. The processed protein and text data are provided as a soft prompt to the LLM. Leveraging the flexible text generation capabilities of the LLM, our model effectively follows residue-level protein instructions, significantly enhancing its utility in protein-related tasks. As shown in Figure 1, the answering effect of our model can provide effective insights for human to explore proteins in a Q&A manner familiar to them.
@@
To evaluate the model’s performance, we conducted comparative experiments on the Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt datasets, and performed zero-shot experiments on a subset of UniProtQA to validate generalization. We used common metrics like BLEU and ROUGE, as well as online KIMI qin2024mooncake scoring metrics and expert evaluations, ensuring the reliability and relevance of the answers. Experimental results show that Prot2Chat which integrates protein sequences and structures using a unified encoder and fuses text information at an early stage outperforms those using sequence-only pretraining, ESM-based encoders lin2022language , and multi-encoders. Besides, its effectiveness far outperforms that of late-fusion, which means that text information is only utilized when the LLM generates answers. These findings highlight the effectiveness of early-fusion strategies for integrating multimodal protein representations. Furthermore, the high consistency between expert evaluations and online KIMI assessments confirms the robustness of Prot2Chat.
@@
    We extended the existing structure encoder ProteinMPNN to realize structure and sequence early-fusion without the need for training.
@@
    Based on this protein encoder, we designed a text-aware protein-text adapter module. By re-utilizing the LLM, we achieved question-based compression of protein information, early-fusion of text and protein information, and implemented a lightweight and efficient protein LLM with only 109 million training parameters.
@@ 2 Related work
2.1 Protein Representation Learning
@@
Proteins are fundamental components of cells, essential for their biological activities and diverse functions. Previous studies on protein characterization have explored various methods to learn protein representations based on different forms of protein information. Protein sequences, often referred to as the ”language of life,” have been extensively studied using advanced natural language processing techniques. For instance, Tranception notin2022tranception employs the Transformer vaswani2017attention model to encode amino acid sequences, capturing relationships between residues and autoregressively reconstructing protein sequences from large-scale databases. Similarly, sequential modeling approaches such as ESM lin2023evolutionary leverage masked language modeling (MLM) to develop attention patterns that correspond to residue-residue contact maps, enhancing sequence-based protein representations. On the other hand, structure-based approaches gligorijevic2021structure directly indicate protein function and encode geometric information of the protein for topology-sensitive tasks such as protein property prediction. Foldseek van2022foldseek introduces the concept of a structural alphabet, encoding protein structures into a discrete representation space. Similarly, Saprot su2023saprot introduces a structure-aware vocabulary, embedding structural information into model input to enhance representational capacity, achieving significant success in protein function prediction tasks.
2.2 Multimodal Alignment
@@
Enabling LLMs to understand extra modalities has been a fast-evolving research area, with notable examples including image-text models li2023blip , video-text models like VideoLlama zhang2023video , audio-text models such as Macaw-LLM lyu2023macaw , and molecular-text models like MolTC fang2024moltc . This line of research was pioneered by advancements in visual language modeling (VLM), to enable LLMs to understand images, leading VLM methods adopt different strategies. Some, like BLIP-2 li2023blip , employ nonlinear and expressive cross-modal projectors, while others, like PaLM-E driess2023palm , utilize visual encoders and fine-tune LLMs on multimodal datasets. These approaches have also been increasingly applied to protein analysis. For instance, Galactica taylor2022galactica leverages scientific literature to model protein sequences and SMILES representations, enabling the model to interpret sequence properties. Some of these methods attempt to adopt a Q&A approach that is closer to human habits to explore and understand the functions of proteins. ProtNote char2024protnote , utilize two encoders to encode protein sequence information and text information, respectively, and Multilayer Perceptron (MLP) to fuse these inputs, achieving the goal of using text to achieve supervised and zero-shot protein functional prediction. ProtChatGPT wang2024protchatgpt uses two independent modules to encode sequence and structural information respectively, and aligns them with natural language to explore protein functions. InstructProtein wang2023instructprotein employs a supervised training framework based on instruction generation using knowledge graphs, enabling bidirectional generation between natural language and protein language. This model can predict protein functional descriptions and generate protein sequences based on natural language prompts. Prot2Text abdine2024prot2text combines graph neural networks (GNNs) with LLMs to predict protein functions in free-text form. FAPM xiang2024fapm utilizes a contrastive learning framework to implement a generative-like Q&A model, while BiomedGPT luo2023biomedgpt bridges the gap between biological and human natural language through a multimodal generative pre-trained transformer (GPT). This allows users to ”communicate” with biological modalities using free-text queries, facilitating interactions with biological data in natural language.
Figure 2: Model Structure of Prot2Chat. The red font represents the input, the snowflake represents freezing, and the flame represents the parameters to be trained. We obtain the embedding with multi-dimensional feature fusion of proteins from protein structure and sequence information through an encoder. Meanwhile, we get the question vector from the question text. Then, we conduct early-fusion and alignment of this vector with protein information to obtain the soft prompt. Finally, we input the soft prompt along with the question text into the LLM to get the answer.
Table 1: Comparison of input-output of protein Q&A models, ”Description” is a text related to protein functions, such as the description of Gene Ontology (GO) terms, ”[Structure+Sequence+Question]” represents the early-fusion of structure, sequence and text.
@@
Prot2Chat [Structure+Sequence+Question]&Question Free Text
Table 2: The count and division of the datasets we used.
@@
Prot2Chat 35.85 57.21 38.09 50.51
Table 4: Results of the ablation experiments in Mol-instructions dataset. The best performance is marked in bold.
Model BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L
Prot2Chat 35.85 57.21 38.09 50.51
@@
Prot2Chat (Zero-shot) 2.23 14.73 1.73 10.10
Prot2Chat (Fine-tuned) 6.72 15.71 9.25 15.57
@@
Model 1st 2nd 3rd 4th Avg.
Prot2Chat 386 242 13 9 1.45
BioMedGPT-LM-10B 26 57 326 241 3.20
LLaMA3-FT with sequence 18 22 252 358 3.46
Evola-10B 220 329 59 42 1.88
@@
Model 1st 2nd 3rd 4th Avg.
Prot2Chat 359 266 17 8 1.49
BioMedGPT-LM-10B 28 58 334 230 3.18
LLaMA3-FT with sequence 11 21 239 379 3.52
Evola-10B 252 305 60 33 1.81
@@
[Uncaptioned image] [O58736]Using the protein structure and the following protein sequence as input, assess its potential function by comparing it to known proteins with similar sequences or structures: The provided protein sequence is approximately 240 amino acids long and shows no significant similarity to known proteins in the database. However, it can be analyzed for potential functional regions and motifs using various bioinformatic tools. Here are some observations and predictions: 1. Transmembrane regions: The sequence contains several regions that are predicted to be transmembrane, including the following: 24-45 amino acids (ILLLVGMPGSGKGEVAKAFRRRGIPVI) 63-82 amino acids (MGGDAIREEAEKRGIP KTPEGLKYVSLKVREE) With the provided structure and amino acid sequence, the protein is expected to display adenylate kinase activity and participate in the phosphorylation, ATP biosynthetic process, phosphorylation. The protein with the structure and amino acid sequence has been analyzed, and its structure suggests a potential involvement in ATP binding.